OpenStudy (anonymous):

Use the given function value, and trigonometric identities (including the cofunction identities), to find the indicated trigonometric functions. sec θ = 5 cot(90° − θ)= sin θ=

OpenStudy (anonymous):

do you know what quadrant you are in?

OpenStudy (anonymous):

does not provide that information

OpenStudy (anonymous):

then you cannot really do the problem, but we can assume we are in quadrant 1 and make everything positive

OpenStudy (anonymous):

first we find \(\cos(x)=\frac{1}{\sec(x)}=\frac{1}{5}\)

OpenStudy (anonymous):

then by pythagoras we find \(\sin(x)\) 

OpenStudy (anonymous):

third side via \(\sqrt{5^2-1^2}=\sqrt{24}\) so \(\sin(x)=\frac{\sqrt{24}}{5}\)

OpenStudy (anonymous):

co-function identity tells us \(\cot(90-x)=\tan(x)\) and since we have cosine and sine, we have tangent, that is \(\tan(x)=\frac{\sqrt{24}}{1}=\sqrt{24}\)